Marigold Company purchased a new plant asset on April 1, 2020, at a cost of $723,000. It was estimated to have a service life of 20 years and a salvage value of $59,400. Marigold’ accounting period is the calendar year.

Compute the depreciation for this asset for 2020 and 2021 using the sum-of-the-years'-digits method. (Round answers to 0 decimal places, e.g. 45,892.)

Depreciation for 2020

$enter a dollar Amountt

Depreciation for 2021

Compute the depreciation for this asset for 2020 and 2021 using the double-declining-balance method. (Round answers to 0 decimal places, e.g. 45,892.)

Depreciation for 2020

$enter a dollar amount

Depreciation for 2021

$enter a dollar amount
Answer #1

Compute the depreciation for this asset for 2020 and 2021 using the sum-of-the-years'-digits method. (Round answers to 0 decimal places, e.g. 45,892.)

Depreciation for 2020

663600*20/210*9/12 = 47400

Depreciation for 2021

663600*20/210*3/12+663600*19/210*9/12 = 60830

Compute the depreciation for this asset for 2020 and 2021 using the double-declining-balance method. (Round answers to 0 decimal places, e.g. 45,892.)

Depreciation for 2020

723000*10%*9/12 = 54225

Depreciation for 2021

723000*10%*3/12+723000*90%*10%*9/12 = 66878